The before-tax income for Flounder Co. for 2017 was $94,000 and $84,300 for 2018. However, the accountant noted that the following errors had been made:

1. Sales for 2017 included amounts of $35,000 which had been received in cash during 2017, but for which the related products were delivered in 2018. Title did not pass to the purchaser until 2018.
2. The inventory on December 31, 2017, was understated by $8,600.
3. The bookkeeper in recording interest expense for both 2017 and 2018 on bonds payable made the following entry on an annual basis.

Interest Expense

15,000

     Cash

15,000

The bonds have a face value of $250,000 and pay a stated interest rate of 6%. They were issued at a discount of $14,000 on January 1, 2017, to yield an effective-interest rate of 7%. (Assume that the effective-yield method should be used.)
4. Ordinary repairs to equipment had been erroneously charged to the Equipment account during 2017 and 2018. Repairs in the amount of $8,200 in 2017 and $9,200 in 2018 were so charged. The company applies a rate of 10% to the balance in the Equipment account at the end of the year in its determination of depreciation charges.


Prepare a schedule showing the determination of corrected income before taxes for 2017 and 2018. (Enter negative amounts using either a negative sign preceding the number e.g. -15,000 or parentheses e.g. (15,000). Round answers to 0 decimal places, e.g. 125.)

2017

2018

Income Before Tax

$

$

Corrections:

Adjustment to Bond Interest ExpenseAdjustment to Bond Interest PayableDepreciation Not Recorded on Capitalized RepairsDepreciation Recorded on Improperly Capitalized RepairsOverstatement of 2017 Ending InventoryRepairs Erroneously Charged to the Equipment AccountRepairs Not Charged to Equipment AccountSales Erroneously Excluded in 2017 IncomeSales Erroneously Included in 2017 IncomeUnderstatement of 2017 Ending Inventory

Adjustment to Bond Interest ExpenseAdjustment to Bond Interest PayableDepreciation Not Recorded on Capitalized RepairsDepreciation Recorded on Improperly Capitalized RepairsOverstatement of 2017 Ending InventoryRepairs Erroneously Charged to the Equipment AccountRepairs Not Charged to Equipment AccountSales Erroneously Excluded in 2017 IncomeSales Erroneously Included in 2017 IncomeUnderstatement of 2017 Ending Inventory

Adjustment to Bond Interest ExpenseAdjustment to Bond Interest PayableDepreciation Not Recorded on Capitalized RepairsDepreciation Recorded on Improperly Capitalized RepairsOverstatement of 2017 Ending InventoryRepairs Erroneously Charged to the Equipment AccountRepairs Not Charged to Equipment AccountSales Erroneously Excluded in 2017 IncomeSales Erroneously Included in 2017 IncomeUnderstatement of 2017 Ending Inventory

Adjustment to Bond Interest ExpenseAdjustment to Bond Interest PayableDepreciation Not Recorded on Capitalized RepairsDepreciation Recorded on Improperly Capitalized RepairsOverstatement of 2017 Ending InventoryRepairs Erroneously Charged to the Equipment AccountRepairs Not Charged to Equipment AccountSales Erroneously Excluded in 2017 IncomeSales Erroneously Included in 2017 IncomeUnderstatement of 2017 Ending Inventory

Adjustment to Bond Interest ExpenseAdjustment to Bond Interest PayableDepreciation Not Recorded on Capitalized RepairsDepreciation Recorded on Improperly Capitalized RepairsOverstatement of 2017 Ending InventoryRepairs Erroneously Charged to the Equipment AccountRepairs Not Charged to Equipment AccountSales Erroneously Excluded in 2017 IncomeSales Erroneously Included in 2017 IncomeUnderstatement of 2017 Ending Inventory

Corrected Income Before Tax

$

$
Answer #1
Schedule to determine correct Income before tax :
S.No. 2017 2018 DETAILS:
Incorrect Income before tax 94000 84300
1 Unearned revenue, earned in 2018 -35000 35000 SALES OF 2017 DECREASES AND OF 2018 INCREASES
2 Understated inventory 8600 -8600 COGS IN 2017 WOULD DECREASE & IN 2018 INCREASES
3 short interest booked -1521 Amortisation in 2017 = (250000-14000)*7% -15000 = 1520 & it forms part of interest
short interest booked -1626 Amortisation in 2018 = (250000-14000+1521)*7% -15000 = 1626
4 Repairs booked -8200 -9200 repairs charged to the Renenues and income reduces accordingly.
Excess depreciation 820 excess depreciation = 8200*10%=820, it should not be charged
Excess depreciation                     1740 excess depreciation = (8200+9200)*10%=1740
Correct Income before tax 58699 101614